A solution containing 2.610 g of cuso45Hzo (molar mass 249.7 g mol-1 is completely reacted with excess Zn metal in a constant pressure calorimeter, according to: Cu Zn Cu Zn 2+. (a) Determine 4H for the reaction as written above if a temperature change of 4.319 K is measured in the reaction mixture. Assume that the combined heat capacity of the calorimeter and solution is 526.37 J Ki. (b) Is the enthalpy change determined in (i) the standard enthalpy change (4H) for the reaction?
Answer #1

a) \large \Delta​H = qp = Cp\large \DeltaT where Cp = 526.37J/K and \large \Delta​T = 4.319 K

= 526.37 J/K x 4.319 K = 2273.4 J

Number of moles of CuSO4.5 H2O = given mass/molar mass = 2.610g/249 g/mol = 0.0105 mol

Molar enthalpy, \large \Delta​H = 2273.4/0.0105 J/mol = 216.5 kJ/mol

b) standard enthalpy change is the enthalpy measured at 1 bar pressure and 298 K temperature. So the enthalpy is not the standard enthlpy change of the reaction.